# Flow is slow at the nanoscale: Revisiting the Green–Kubo relation for friction

### v1

@@ III.2 Connection to the hydrodynamic slip boundary
While the GK relation provides a microscopic expression for the effective friction, more often than not, one is interested in the intrinsic friction or the slip length of a particular interface. Since the constitutive relation of the effective friction has been established, it is relatively straightforward to obtain a relationship between the effective friction and the slip length from macroscopic hydrodynamics. Specifically, we can solve the Navier–Stokes equation[59, 60] subject to partial slip boundary conditions like Eq. 1 at the solid wall to obtain the Poiseuille flow profile of the fluid. From the solution, we then obtain the fluid velocity as well as its frictional force per unit area, the ratio of which gives an expression for the effective friction coefficient. Here, we will simply state the results for the common cases of 2D and 1D confinement, with details presented in the SI.
@@
from an EMD simulation with a constraint keeping the liquid linear momentum zero in the in-plane $(x,y)$ directions, where $\bm{\mathit{F}}_{\parallel}=(F_{x},F_{y})$ denotes the in-plane lateral force on the liquid.
For 1D confinement, we consider Poiseuille flow of a fluid through a cylindrical tube of radius $R$. In an analogous manner, the effective friction coefficient can be obtained from classical hydrodynamics to be
@@
from an EMD simulation with a constraint to keep the liquid momentum zero along the axial direction ($z$) of the tube, where $F_{z}$ denotes the force on the liquid in the axial direction.
@@
The pair of Eqs. 32/33 and 35/36 provides a way to obtain hydrodynamic slippage directly from fluctuating forces in microscopic equilibrium simulations. Once the slip length $b$ is obtained for a particular surface, the intrinsic friction coefficient $\lambda_{\mathrm{intr}}$ follows directly from Eq. 3. In Fig. 2(b), we validate that $\lambda_{\mathrm{intr}}$ obtained by this method (GK+CHT) are in excellent agreement with reference NEMD simulations for surfaces of different wettability. In comparison, the BB expression in Eq. 4 significantly underpredicts the friction for surfaces with higher wettability.
@@
In writing Eqs. 31 and 34, we have made a simplifying approximation when treating the liquid as a continuum in CHT that the hydrodynamic boundary applies at the first plane of solid atoms in contact with the liquid. Therefore, the solution for the velocity profile from the Navier–Stokes equation spans across the total height of the channel or the diameter of the cylindrical tube. However, microscopically, there is potentially an offset in the location of the hydrodynamic boundary due to the excluded volume at each liquid–solid interface. In the SI, we assess the sensitivity of $\lambda_{\mathrm{eff}}$ to such an offset, noting that care should be taken in placing the hydrodynamic boundary as the channel height decreases and/or the surfaces becomes less slippery as $\lambda_{\mathrm{eff}}$ is most sensitive in the limit $H\to 0$ and $b\to 0$. Generally speaking, in cases where $H>1\,\mathrm{nm}$, placing the hydrodynamic boundary at the first atomic plane of the solid is a reasonable approximation.
@@
It is important to note that the appropriate type of flow used to give the mapping between CHT and the effective friction is Poiseuille flow, where the fluid has been driven by a pressure gradient or a body force. This is because under confinement, when the liquid is in contact with two separate solid substrates, the solid’s frame of reference is only well-defined when both substrates have the same velocity as each other. Meanwhile, in the case of Couette flow, the two solid substrates are moving at different velocities in order to shear the fluid. However, as the slip length $b$ is an intrinsic property, once its value is known, predictions of the velocity profile can be obtained from CHT. For example, we show the predictions for the symmetric $\alpha=3$ channel, both for Couette flow and Poiseuille flow in Fig. 2(c), and for other cases extensively in the SI. In all cases, we find very good agreement with reference NEMD simulations.
@@ IV Application: Water friction under confinement
Figure 3: Water friction under confinement. (a) The curvature dependence of the slip length $b$ for water on different surfaces including carbon nanotubes (CNTs), boron nitride nanotubes (BNNTs), nanotubes made of Lennard–Jones particles (LJNTs), flat graphene sheet (GRA) and flat hexagonal boron nitride sheet (hBN) obtained from simulations. The dashed lines are fits to simulation data point. The effective friction under 1D and 2D confinement is shown in (b) and (c) as a function of the tube radius $R$ and channel height $H$. We consider the same surfaces in (a), with addition of a solid substrate of Lennard–Jones particles of high wettability with $\alpha=3$ (LJ). The data points are obtained from simulations and the lines show predictions from classical hydrodynamics theory (CHT). The dash lines show the results using no-slip boundary conditions while solid lines account for finite slippage at the interface. The shaded regions indicate where CHT begins to break down.
@@
Historically, the no-slip boundary condition $b=0$ can successfully describe much of everyday phenomena involving fluid dynamics [61, 62]. For water flowing through macroscopic tubes and channels, it provides a very good approximation, regardless of the material of the solid wall. At these larger scales, the effective friction increases monotonically as the tube becomes smaller $\lambda^{\mathrm{1D}}_{\rm eff}\sim 4\eta/R$ or as the channel height decreases $\lambda^{\mathrm{2D}}_{\rm eff}\sim 12\eta/H$. It has also long been established that as the length scale of confinement keeps shrinking in size, the no-slip boundary condition is expected to break down [63, 64] as a larger proportion of the confined fluid can feel the interface. Thus, the first question we will address is “At which confinement length scale do surface effects begin to manifest as a deviation from the no-slip boundary approximation?” The answer will depend on various factors, including the material and the curvature of confinement.
@@
To answer this question, we will consider 1D confinement in single-walled carbon and boron nitride nanotubes of different radii and 2D confinement in channels of different heights formed by two graphene or boron nitride sheets. For the interatomic interaction, we employ a machine-learned potential where the electronic structure is accurately trained at the level of density functional theory [17]. While we consider these low-dimensional materials for their relevance to applications in nanofluidics, we stress the generality of the results obtained to confinement by other solid surfaces of other liquids. While the microscopic mechanisms giving rise to friction at the interface are very diverse (corrugation [17, 14], phonon coupling [13], quantum friction [65, 20], defects [47, 46, 42], etc.) and we expect them to depend sensitively on the molecular details of the interface, the macroscopic hydrodynamic behavior would be principally governed by the slip length. Therefore, in addition to these realistic surfaces, we also consider water confined in Lennard–Jones nanotubes and between walls of different wettability.
@@
For each surface, we first obtain the slip length from EMD simulations using the pair of Eqs. 32/33 and 35/36. Between the different wall materials, the graphene surface has the highest slip length ($b=16\,$nm), followed by boron nitride ($b=3.6\,$nm) and the Lennard–Jones surface of high wettability where there is essentially no slippage. In the case of 1D confinement, in addition to surface effects due to different confining materials, curvature effects also come into play since curving up a surface can change its free energy surface significantly. We show the curvature-dependent slip length $b_{R}$ for three different surfaces in Fig. 3(a). In all cases, curving up the surface to form nanotubes leads to a decrease in the corrugation of the surface, reflected by an increase in $b_{R}$ as $R$ decreases. This enhancement is most prominent for tubes of $R\lesssim 5\,$nm with the slippery carbon surface. Our values are broadly in agreement with previous simulations and experiments of water on graphene and boron nitride surfaces [14, 16, 10, 12, 17, 20].
@@
To assess the overall effect of changing the radius on water flow in 1D confinement, using $b_{R}(R)$ obtained from a non-linear fit of the simulation data in Eq. 31, we show in Fig. 3(b) the effective friction $\lambda^{\mathrm{1D}}_{\mathrm{eff}}$ as a function of $R$. There are two competing effects dictating the behavior of $\lambda^{\mathrm{1D}}_{\mathrm{eff}}$. For tubes with diameters on the micrometer scale and beyond, $\lambda^{\mathrm{1D}}_{\mathrm{eff}}$ increases as $R$ decreases for all surfaces, asymptotic to the no-slip prediction. As $R$ reaches the nanometric regime, since it is now comparable to $b_{R}$, the effective friction in the fluid is no longer agnostic to the amount of slippage taking place at the wall. As a result, the enhancement in $b_{R}$ due to curvature effects starts to dominate, decreasing $\lambda^{\mathrm{1D}}_{\mathrm{eff}}$ as $R$ decreases. Therefore, a crossover in the dependence of $\lambda^{\mathrm{1D}}_{\rm eff}$ on $R$ can be observed at $R\sim 10\,$nm and is most significant for surfaces of the highest slippage. We note that experiments [10] with multi-walled carbon nanotubes have reported a radius dependence of slip length up to 505050\,50nm, so the length scale of the crossover could be as large as tens of nm. For 2D confinement, since there is no curvature effect, any surface effect would be solely due to interactions between the solid and the fluid. In this case, instead of a crossover, a deviation to the no-slip prediction is seen for the asymptotic behavior of $\lambda^{\mathrm{2D}}_{\mathrm{eff}}$ as the channel height $H$ decreases, as shown in Fig. 3(c) for water confined in symmetric channel formed with graphene, hexagonal boron nitride and the Lennard–Jones solid substrate. Similarly, the length scale of this deviation is in the tens of nm and is most significant for surfaces of the highest slippage.
@@
To answer this question, we compared the CHT predictions to $\lambda^{\mathrm{2D}}_{\rm eff}$ computed from EMD simulations for various channel heights down to the sub-nanometer regime, shown as data points in Fig. 3(c). We found that in all confining materials considered, CHT predictions hold remarkably well down to $H\sim 1\,$nm, where the channels can accommodate only three layers of water. The deviation between CHT and simulations is only significant for smaller channels with a bilayer and monolayer of water, where we found that $\lambda^{\mathrm{2D}}_{\rm eff}$ from simulations is extremely sensitive to changes in the density and interfacial structure of the fluid. The breakdown of CHT below 111\,1nm is perhaps not surprising since with fewer than three layers of water, one can no longer sensibly define a continuum bulk region. In fact, the length scale of 111\,1nm has been suggested both experimentally [66, 67] and theoretically [68, 69] as the limit for the validity of the notion of a bulk shear viscosity. Nevertheless, the robustness of a continuum theory to describe water flow down to such a small scale is extremely satisfying. It is also reminiscent of the validity of other macroscopic theories, e.g. continuum mechanics [70], the Kelvin equation [71] or dielectric continuum theory[72, 73, 74, 75], down to the nanoscale.
@@ V Conclusion
In this work, through the use of the projection operator formalism and linear response theory, we have derived a GK relation for the liquid–solid friction by considering the stochastic motion of a liquid droplet on a solid surface. The force autocorrelation function in the expression is sampled using constrained dynamics in which the momentum of the liquid is conserved. Not only does such an expression help resolve the long-standing plateau problem associated with previous studies of friction using molecular simulations, it also sheds light on the physical meaning of the obtained friction coefficient: $\lambda_{\rm eff}$ measures an effective friction and not the intrinsic interfacial friction. Our expression can be applied generally to finite systems with surfaces of very low or very high friction. In contrast, the widely used expression by Bocquet and Barrat was derived originally for a semi-infinite system[30, 26]. When applied to finite systems, in the context of our work, it amounts to a reasonable approximation only when the slip length is much greater than the confining length scale.
By linking the microscopic expression for friction to macroscopic hydrodynamics, we show how the hydrodynamic slip length can be obtained faithfully, both for 1D and 2D fluid flow. Going forwards, we believe that such a connection is important to understand the flow of liquids across length scales on more complex surfaces where slippage is determined by the interplay of the material’s electronic properties, wettability, surface charge and the presence of defects. This work also lays the foundation for understanding interfacial effects in electrokinetic phenomena and other transport properties under confinement [76, 77].
Finally, as an application of our approach, we assess the ability of continuum hydrodynamics in describing water flow at the nanoscale. We show that the no-slip boundary condition often employed at the macroscale begins to break down for confinement lengths below 10–100 nm. However, once finite slippage is accounted for using partial slip boundary conditions, continuum hydrodynamics remains remarkably robust down to 1 nm confinement. Our results also emphasize an obvious implication for physics at the nanoscale: As a fluid becomes increasingly confined, the relative importance of effects at the interface are enhanced relative to those in bulk. If one aims to exploit flow-induced effects that are inherently interfacial, such as generating pressure-driven electrical currents, [78, 1] there is potential gain to nanoscale confinement. However, if one is simply interested in minimizing frictional effects to maximize the total flux per unit driving force, then this work makes clear that “bigger is better”.

### v2

@@ III.2 Connection to the hydrodynamic slip boundary
While the GK relation provides a microscopic expression for the effective friction, more often than not, one is interested in the intrinsic friction or the slip length of a particular interface. Since the constitutive relation of the effective friction has been established, it is relatively straightforward to obtain a relationship between the effective friction and the slip length from macroscopic hydrodynamics. Specifically, we can solve the Navier–Stokes equation[60, 61] subject to partial slip boundary conditions like Eq. 1 at the solid wall to obtain the Poiseuille flow profile of the fluid. From the solution, we then obtain the fluid velocity as well as its frictional force per unit area, the ratio of which gives an expression for the effective friction coefficient. Here, we will simply state the results for the common cases of 2D and 1D confinement, with details presented in the supplementary material.
@@
from an EMD simulation with a constraint keeping the liquid’s linear momentum zero in the in-plane $(x,y)$ directions, where $\bm{\mathit{F}}_{\parallel}=(F_{x},F_{y})$ denotes the in-plane lateral force on the liquid.
For 1D confinement, we consider Poiseuille flow of a fluid through a cylindrical tube of radius $R$. In an analogous manner, the effective friction coefficient can be obtained from classical hydrodynamics as
@@
from an EMD simulation with a constraint keeping the liquid’s momentum zero along the axial direction ($z$) of the tube, where $F_{z}$ denotes the force on the liquid in the axial direction.
@@
The pair of Eqs. 31/32 and 34/35 provides a way to obtain hydrodynamic slippage directly from fluctuating forces in microscopic EMD simulations. In Fig. 2(b), we validate that the slip lengths $b$ obtained by this method (labeled GK*+CHT) are in excellent agreement with results from reference NEMD simulations for surfaces of different wettability. (To obtain $b$ from NEMD, we have measured $\lambda_{\rm eff}$ and used the result in the CHT expression, rather than attempt to fit the flow profiles directly.) In contrast, using the BB expression in Eq. 4 combined with Eq. 3, as is typically done in the literature, significantly overpredicts the slip length as the wettability increases.
@@
While sources of uncertainty in the effective friction measured with the GK relation arise from statistical error, the value of slip length is also impacted by the choice of hydrodynamic boundary position $z_{0}$ (which enters Eqs. 31 and 34 implicitly through $H$ and $R$). While approaches exist for determining $z_{0}$, see, e.g., Refs. 62 and 63, here we make the pragmatic choice to place $z_{0}$ at the first plane of solid atoms in contact with the liquid. With this choice, the solution for the velocity profile from the Navier–Stokes equation spans the total height of the channel or the diameter of the cylindrical tube. However, microscopically, there is potentially an offset in the location of the hydrodynamic boundary due to the excluded volume at each liquid–solid interface. In the supplementary material, we assess the sensitivity of our results to such an offset. We find that, when the fluid’s friction is dominated by viscous dissipation (i.e., in tubes or channels that are large compared to the slip length), results are robust to reasonable choices of $z_{0}$. As $H$ or $R$ tends to zero, obtained values of $b$ become increasingly sensitive to the choice of $z_{0}$; generally speaking, in cases where $H>1\,\mathrm{nm}$, placing the hydrodynamic boundary at the first atomic plane of the solid is a reasonable approximation.
@@
It is important to note that the appropriate type of flow used to give the mapping between CHT and the effective friction is Poiseuille flow, where the fluid has been driven by a pressure gradient or a body force. This is because under confinement, when the liquid is in contact with two separate solid substrates, the solid’s frame of reference is only well-defined when both substrates have the same velocity as each other. Note that the mapping to Poiseuille flow enables calculation of friction in cylindrical geometries. Meanwhile, in the case of Couette flow, the two solid substrates are moving at different velocities in order to shear the fluid. However, as the slip length $b$ is an intrinsic property, once its value is known, predictions of the velocity profile can be obtained from CHT. For example, we show the predictions for the symmetric $\alpha=3$ channel, both for Couette flow and Poiseuille flow in Fig. 2(c), and for other cases extensively in the supplementary material. In all cases, we find very good agreement with reference NEMD simulations.
@@ IV Application: Water friction under confinement
Figure 3: Water friction under confinement. (a) The curvature dependence of the slip length $b$ for water on different surfaces including carbon nanotubes (CNTs), boron nitride nanotubes (BNNTs), nanotubes made of Lennard–Jones particles (LJNTs), a flat graphene sheet (GRA) and a flat hexagonal boron nitride sheet (hBN) obtained from simulations. The effective friction under 1D and 2D confinement is shown in (b) and (c) as a function of the tube radius $R$ and channel height $H$. The data points are obtained by the GK relation using constrained EMD simulations. For 2D confinement, the solid lines are predictions from CHT (Eq. 30) parameterized on the slip length of the largest channel. For 1D confinement, the dashed lines are from Eq. 33 using $b_{R}(R)$ fitted to the simulation data. The dotted lines show the results using no-slip boundary conditions while solid lines account for finite slippage at the interface. The shaded regions indicate where CHT begins to break down.
@@
Historically, the no-slip boundary condition $b=0$ can successfully describe much of everyday phenomena involving fluid dynamics [64, 65]. For water flowing through macroscopic tubes and channels, it provides a very good approximation, regardless of the material of the solid wall. At these larger scales, the effective friction increases monotonically as the tube becomes smaller $\lambda^{\mathrm{1D}}_{\rm eff}\sim 4\eta/R$ or as the channel height decreases $\lambda^{\mathrm{2D}}_{\rm eff}\sim 12\eta/H$. It has also long been established that as the length scale of confinement keeps shrinking in size, finite slippage at the interface can no longer be ignored [66, 67] as a larger proportion of the confined fluid can feel the interface. Thus, the first question we will address is “At which confinement length scale do surface effects begin to manifest as a deviation from the no-slip boundary approximation?” The answer will depend on various factors, including the material and the curvature of confinement.
@@
To answer this question, we will consider 1D confinement in single-walled carbon and boron nitride nanotubes of different radii and 2D confinement in channels of different heights formed by two graphene or boron nitride sheets. For the interatomic interaction, we employ a machine-learned potential where the electronic structure is accurately trained at the level of density functional theory [17]; the walls are flexible in these simulations. While we consider these low-dimensional materials for their relevance to applications in nanofluidics, we stress the generality of the results obtained to confinement by other solid surfaces of other liquids. While the microscopic mechanisms giving rise to friction at the interface are very diverse (corrugation [17, 14], phonon coupling [13], quantum friction [68, 20], defects [47, 46, 42], etc.) and we expect them to depend sensitively on the molecular details of the interface, the macroscopic hydrodynamic behavior would be principally governed by the slip length. Therefore, in addition to these realistic surfaces, we also consider water confined in Lennard–Jones nanotubes and between walls of different wettability.
@@
For each surface, we first obtain the slip length from EMD simulations using the pair of Eqs. 31/32 and 34/35. Between the different wall materials, the graphene surface has the highest slip length ($b=16\,$nm), followed by boron nitride ($b=3.6\,$nm) and the Lennard–Jones surface of high wettability where there is essentially no slippage. In the case of 1D confinement, in addition to surface effects due to different confining materials, curvature effects also come into play since curving up a surface can change its free energy surface significantly. We show the curvature-dependent slip length $b_{R}$ for three different surfaces in Fig. 3(a). In all cases, curving up the surface to form nanotubes leads to a decrease in the corrugation of the surface, reflected by an increase in $b_{R}$ as $R$ decreases. This enhancement is most prominent for tubes of $R\lesssim 5\,$nm with the slippery carbon surface. Our values are broadly in agreement with previous simulations and experiments of water on graphene and boron nitride surfaces [14, 16, 10, 12, 17, 20].
@@
To assess the overall effect of changing the radius on water flow in 1D confinement, using $b_{R}(R)$ obtained from a non-linear fit of the simulation data in Eq. 30, we show in Fig. 3(b) the effective friction $\lambda^{\mathrm{1D}}_{\mathrm{eff}}$ as a function of $R$. There are two competing effects dictating the behavior of $\lambda^{\mathrm{1D}}_{\mathrm{eff}}$. For tubes with diameters on the micrometer scale and beyond, $\lambda^{\mathrm{1D}}_{\mathrm{eff}}$ increases as $R$ decreases for all surfaces, asymptotic to the no-slip prediction. As $R$ reaches the nanometric regime, since it is now comparable to $b_{R}$, the effective friction in the fluid is no longer agnostic to the amount of slippage taking place at the wall. As a result, the enhancement in $b_{R}$ due to curvature effects starts to dominate, decreasing $\lambda^{\mathrm{1D}}_{\mathrm{eff}}$ as $R$ decreases. Therefore, a crossover in the dependence of $\lambda^{\mathrm{1D}}_{\rm eff}$ on $R$ can be observed at $R\sim 10\,$nm and is most significant for surfaces of the highest slippage. We note that experiments [10] with multi-walled carbon nanotubes have reported a radius dependence of slip length up to 505050\,50nm, so the length scale of the crossover could be as large as tens of nm. For 2D confinement, since there is no curvature effect, any surface effect would be solely due to interactions between the solid and the fluid. In this case, instead of a crossover, a deviation to the no-slip prediction is seen for the asymptotic behavior of $\lambda^{\mathrm{2D}}_{\mathrm{eff}}$ as the channel height $H$ decreases, as shown in Fig. 3(c) for water confined in symmetric channels of graphene, hexagonal boron nitride and the Lennard–Jones solid substrates. Similarly, the length scale of this deviation is in the tens of nm and is most significant for surfaces of the highest slippage.
@@
Since the GK relation we have derived is independent of CHT, we can parameterize CHT with the slip length obtained from a particular geometry, and then assess how well this parameterized CHT performs as we decrease the confinement length. To this end, we compare the effective friction predicted by CHT (Eq. 30) with that computed from the GK relation for various channel heights down to the sub-nanometer regime, shown as data points in Fig. 3(c). We find that in all confining materials considered, CHT predictions hold remarkably well down to $H\sim 1\,$nm, where the channels can accommodate only three layers of water. The deviation between CHT and simulations is only significant for smaller channels with a bilayer and monolayer of water, where we find that $\lambda^{\mathrm{2D}}_{\rm eff}$ from simulations is extremely sensitive to changes in the density and interfacial structure of the fluid. The breakdown of CHT below 111\,1nm is perhaps not surprising since with fewer than three layers of water, one can no longer sensibly define a continuum bulk region. In fact, the length scale of 111\,1nm has been suggested both experimentally [69, 70] and theoretically [71, 72] as the limit for the validity of the notion of a bulk shear viscosity. Nevertheless, the robustness of a continuum theory to describe water flow down to such a small scale is extremely satisfying. It is also reminiscent of the validity of other macroscopic theories, e.g., continuum mechanics [73], the Kelvin equation [74] or dielectric continuum theory[75, 76, 77, 78], down to the nanoscale.
@@ V Conclusion
In this work, through the use of the projection operator formalism and linear response theory, we have derived a GK relation for the liquid–solid friction by considering the stochastic motion of a liquid droplet on a solid surface. The force autocorrelation function in the expression is sampled using constrained dynamics in which the momentum of the liquid is conserved. Not only does such an expression help resolve the long-standing plateau problem associated with previous studies of friction using molecular simulations, but it also sheds light on the physical meaning of the obtained friction coefficient: $\lambda_{\rm eff}$ measures an effective friction and not the intrinsic interfacial friction. Our expression can be applied generally to finite systems with surfaces of very low or very high friction. In contrast, the widely used expression by Bocquet and Barrat was derived originally for a semi-infinite system[30, 26]. When applied to finite systems, in the context of our work, it amounts to a reasonable approximation only when the slip length is much greater than the confining length scale.
By linking the microscopic expression for friction to macroscopic hydrodynamics, we show how the hydrodynamic slip length can be obtained faithfully, both for 1D and 2D fluid flow. Going forwards, we believe that such a connection is important to understand the flow of liquids across length scales on more complex surfaces where slippage is determined by the interplay of the material’s electronic properties, wettability, surface charge and the presence of defects. This work also lays the foundation for understanding interfacial effects in electrokinetic phenomena and other transport properties under confinement [79, 80]. We believe our work can provide insights for the development of microscopic descriptions of hydrodynamics based on dynamical density functional theory[81, 63, 82, 83], the non-equilibrium extension of equilibrium classical density functional theory[84, 52, 85], in which the fluid flow is described with mass and momentum density fields.
Finally, as an application of our approach, we assess the ability of continuum hydrodynamics in describing water flow at the nanoscale. While for macroscopic confining geometries assuming no-slip boundary conditions is a reasonable approximation, the impact of finite slippage cannot be ignored for confinement length scales below 10–100 nm. Nevertheless, our results show that, provided the finite slip length is taken into account, continuum hydrodynamics remains remarkably robust down to $\sim$1 nm confinement.
@@
Supplementary Material
Supplementary material includes simulations details and additional supporting results.